# Endpoint Variation and jump inequalities for rough singular integrals

### v1

@@ 1. Introduction
$T_{\Omega}f(x)=\lim_{\epsilon\to 0}T_{\Omega,\epsilon}f(x),\quad T_{\Omega}^{*}f(x)=\sup_{\epsilon>0}|T_{\Omega,\epsilon}f(x)|$
@@
$N_{\lambda}(F_{t}(x):t\in\mathcal{I})=\sup\Big\{N\in{\mathbb{N}}\cup\{0\}:$ $\exists t_{1},t_{2},\cdots,t_{N}\in\mathcal{I},\;0<t_{1}<t_{2}<\cdots<t_{N}\;\text{such that}$
$|F_{t_{i+1}}(x)-F_{t_{i}}(x)|>\lambda\Big\}.$
@@
By a standard limiting argument, Theorem 1.1 recovers the weak $(1,1)$-boundedness of $T_{\Omega}^{*}$ proved recently by Lai [Lai2025].
@@
Let $f\in L^{1}({\mathbb{R}}^{d})$ and $\lambda>0$. By the Calderón-Zygmund decomposition, applied to $f$ at height $\lambda$, there exists a family $\mathcal{Q}$ of disjoint dyadic cubes such that
@@
$\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(S_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}$ $\lesssim V_{1}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(S_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)$
@@
$\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}\lesssim$ $~\mathcal{S}_{2}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)$
@@
where $\mathcal{S}_{2}$ is the short jump operator defined as $\mathcal{S}_{2}(F_{\epsilon}:\epsilon>0)(x)=\left(\sum\limits_{k\in{\mathbb{Z}}}|V_{2}(F_{\epsilon}:\epsilon\in[2^{k-1},2^{k}])(x)|^{2}\right)^{\frac{1}{2}}$. The second term in the above expression is known as the long jump function.
@@ 2.1. Estimate for short jumps
$\left|\left\{x\in E^{c}:\sum_{s\geq 5}\left(\sum\limits_{k\in{\mathbb{Z}}}|V_{2}((H^{s}_{k+j}\chi_{|\cdot|>\epsilon})*B^{k+j-s}(x):\epsilon\in[2^{k-1},2^{k}])|^{2}\right)^{\frac{1}{2}}>\alpha\right\}\right|\lesssim\frac{\|f\|_{1}}{\alpha},\quad\text{for }j=0,1.$
We will prove the inequality for $j=0$, the other estimate is similar. By Chebyshev’s and triangle inequality, it is enough to prove the following for $k\in{\mathbb{Z}}$ and $s\geq 5$,
$\left\|V_{2}((H_{k}^{s}\chi_{|\cdot|>\epsilon})*B^{k-s}:\epsilon\in[2^{k-1},2^{k}])\right\|_{2}^{2}\lesssim 2^{-\delta^{\prime}s}\alpha\sum_{Q\in\mathcal{Q}:|Q|=2^{(k-s)n}}\|b_{Q}\|_{1},\;\text{for some}\;\delta^{\prime}>0.$
Recalling the definition of $V_{2}$, the above inequality is equivalent to
@@
$\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{-s}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}$ $\lesssim\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,in}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}+\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,bd}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}$
@@ 2.2. Estimate for long jumps
$\lesssim$ $\lambda\Big[N_{\frac{\lambda}{2}}\Big(\sum_{s\geq 5}\sum_{j\geq k}H_{j}^{s}*B^{j-s}:k\in{\mathbb{Z}}\Big)(x)\Big]^{\frac{1}{2}}+\lambda\Big[N_{\frac{\lambda}{2}}\Big(\sum_{s\geq 5}(H_{k}^{s}\chi_{2^{k-1}\leq|.|\leq 2^{k}})*B^{k-s}:k\in{\mathbb{Z}}\Big)(x)\Big]^{\frac{1}{2}}$
@@ 2.2.2. Estimate for $L_{1}$:
$\left|\left\{V_{2}\left(\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m}_{u}(x):k\in{\mathbb{Z}}\right)>{c_{1}3^{-d}2^{-\delta_{1}s}\alpha}\right\}\right|,\text{for a fixed}\;u\in\left\{0,\frac{1}{3},\frac{2}{3}\right\}^{d}.$
The advantage of the above expression is that $\operatorname{supp}(H_{j}^{s}*B^{j-s}_{u})\subset\widetilde{K}_{j,m}$, where $\widetilde{K}_{j,m}\in\mathcal{D}_{u}$ and any such two $\widetilde{K}_{j,m}$ and $\widetilde{K}_{j^{\prime},m^{\prime}}$ are either disjoint or one is contained in another. We will now discard the points where there is a large interaction on the scales $j$. This leaves us with points having a controlled overlap in the scales with respect to the parameter $s$. To do so, we initially set
$\mathcal{C}_{0}^{u}=\{(j,m):\exists Q\in\mathcal{Q},\\ |Q|=2^{(j-s)d},\;Q\subseteq K_{j,m}\;\text{and}\;\widetilde{K}_{j,m}\in\mathcal{D}_{u}\},\quad\text{and}\quad\mathcal{X}^{u}=\bigcup\limits_{(j,m)\in\mathcal{C}^{u}_{0}}\widetilde{K}_{j,m}.$
@@
We note that for any $(j,m)\in\mathcal{C}^{u}_{0}$, there exists atleast one cube $Q_{j.m}\in\mathcal{Q}_{j,m}$ with $|Q_{j,m}|=2^{(j-s)d}$ and $Q_{j,m}\subset K_{j,m}$. Thus, by Chebyshev’s inequality, we have
@@
$\sum_{j\geq k}H_{j}^{s}*B^{j-s}_{u}(x)=\sum_{(j,m)\in\mathcal{C}^{u},\;j\geq k}H_{j}^{s}*B^{j-s,m}_{u}(x).$
The advantage of working with the collection $\mathcal{C}^{u}$ is that for any fixed $(j,m)\in\mathcal{C}^{u}$, there are atmost $2^{s(d+\epsilon_{0})}$ indices $(j^{\prime},m^{\prime})\in\mathcal{C}^{u}$ such that $\widetilde{K}_{j,m}\subset\widetilde{K}_{j^{\prime},m^{\prime}}$. Based on this observation, we subdivide the collection $\mathcal{C}^{u}$ as follows. For any $(j,m)\in\mathcal{C}^{u}$, we say $(j,m)\in\mathcal{C}^{u}_{1}$ if there is no other index $(j^{\prime},m^{\prime})\in\mathcal{C}^{u}$ such that $\widetilde{K}_{j^{\prime},m^{\prime}}\subseteq\widetilde{K}_{j,m}$. Inductively, we define $\mathcal{C}^{u}_{l}$ to be the collection of all indices $(j,m)\in\mathcal{C}^{u}\setminus\cup_{i=1}^{l-1}\mathcal{C}^{u}_{i}$ such that there does not exist any index $(j,m)\in\mathcal{C}^{u}\setminus\cup_{i=1}^{l-1}\mathcal{C}^{i}_{v}$ such that $\widetilde{K}_{j^{\prime},m^{\prime}}\subset\widetilde{K}_{j,m}$. By the observation before, the process terminates at the step $l=2^{s(d+\epsilon_{0})}$ and we have $\mathcal{C}^{u}=\bigcup\limits_{l=1}^{2^{s(d+\epsilon_{0})}}\mathcal{C}^{u}_{l}$. Thus, we have that
@@
$\lesssim\frac{2^{2\delta_{1}s}}{\alpha^{2}}\left\|V_{2}\left(\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m}_{u}:k\in{\mathbb{Z}}\right)\right\|_{2}^{2}$
@@ A.1. Estimates of Seeger:
$\int|(A+B)(x)|^{2}dx$ $=\left(\int_{|B(\cdot)|\leq 2^{\rho s}\alpha}+\int_{|A(\cdot)|>2^{\rho s-1}\alpha}+\int_{|(A+B)(\cdot)|>2^{\rho s-1}\alpha}\right)|(A+B)(x)|^{2}dx$
$=:I+II+III$
@@
$\leq$ $\frac{1}{|R|}\int_{R}\left|\sum_{2^{j}\geq\ell(R)}\nu_{j}*|B^{j-s}|(x)-\frac{1}{|R|}\int_{R}\sum_{2^{j}\geq\ell(R)}\nu_{j}*|B^{j-s}|(y)dy\right|dx+\frac{1}{|R|}\int_{R}\left|\sum_{2^{j}<\ell(R)}\nu_{j}*|B^{j-s}|(x)\right|dx$
@@ Acknowledgement
Ankit Bhojak is supported by the Science and Engineering Research Board, Department of Science and Technology, Govt. of India, under the scheme National Post-Doctoral Fellowship, file no. PDF/2023/000708.

### v2

@@ Abstract.
Key words and phrases:
Variation inequalities, jump inequalities, weak type estimate, maximal operator
@@ 1. Introduction
$T_{\Omega}f(x)=\lim_{\epsilon\to 0}T_{\Omega,\epsilon}f(x),\quad T_{\Omega}^{*}f(x)=\sup_{\epsilon>0}|T_{\Omega,\epsilon}f(x)|,$
@@
[…]
$\text{such that}|F_{t_{i+1}}(x)-F_{t_{i}}(x)|>\lambda\Big\}.$
@@
By a standard limiting argument, Theorem 1.1 recovers the weak $(1,1)$ boundedness of $T_{\Omega}^{*}$ proved recently by Lai [Lai2025].
@@
Let $f\in L^{1}({\mathbb{R}}^{d})$ and $\alpha>0$. By the Calderón-Zygmund decomposition, applied to $f$ at height $\alpha$, there exists a family $\mathcal{Q}$ of disjoint dyadic cubes such that
@@
$\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(S_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}$
$\lesssim$ $V_{1}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(S_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)$
@@
$\lambda\left[N_{\frac{\lambda}{2}}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)\right]^{\frac{1}{2}}$
$\lesssim$ $~\mathcal{S}_{2}\left(\sum_{s\geq 5}\sum_{j\in{\mathbb{Z}}}(H_{j}^{s}\chi_{|\cdot|>\epsilon})*B^{j-s}(x):\epsilon\in(0,\infty)\right)$
@@
where $\mathcal{S}_{2}$ is the short jump operator defined as
$\mathcal{S}_{2}(F_{\epsilon}:\epsilon>0)(x)=\left(\sum\limits_{k\in{\mathbb{Z}}}|V_{2}(F_{\epsilon}:\epsilon\in[2^{k-1},2^{k}])(x)|^{2}\right)^{\frac{1}{2}}.$
The second term in the expression (2) is known as the long jump function.
@@ 2.1. Estimate for short jumps
$\left|\left\{x\in E^{c}:\sum_{s\geq 5}\left(\sum\limits_{k\in{\mathbb{Z}}}|V_{2}((H^{s}_{k+j}\chi_{|\cdot|>\epsilon})*B^{k+j-s}(x):\epsilon\in[2^{k-1},2^{k}])|^{2}\right)^{\frac{1}{2}}>\alpha\right\}\right|\lesssim\frac{\|f\|_{1}}{\alpha},$
for $j=0,1$. We will prove the inequality for $j=0$, the other estimate is similar. By Chebyshev’s and triangle inequality, it is enough to prove the following for $k\in{\mathbb{Z}}$ and $s\geq 5$,
$\left\|V_{2}((H_{k}^{s}\chi_{|\cdot|>\epsilon})*B^{k-s}:\epsilon\in[2^{k-1},2^{k}])\right\|_{2}^{2}\lesssim 2^{-\delta^{\prime}s}\alpha\sum_{Q\in\mathcal{Q}:|Q|=2^{(k-s)n}}\|b_{Q}\|_{1},$
for some $\delta^{\prime}>0$. Recalling the definition of $V_{2}$, the above inequality is equivalent to
@@
$\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{-s}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}$
$\lesssim$ $\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,in}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}+\left\|\left(\sum_{i=1}^{N^{k}(x)}|H^{s}_{k,E_{i}^{k}(x)}*B^{k-s}_{i,bd}(x)|^{2}\right)^{\frac{1}{2}}\right\|_{2}^{2}$
@@ 2.2. Estimate for long jumps
$\lesssim$ $\lambda\Big[N_{\frac{\lambda}{2}}\Big(\sum_{s\geq 5}\sum_{j\geq k}H_{j}^{s}*B^{j-s}:k\in{\mathbb{Z}}\Big)(x)\Big]^{\frac{1}{2}}$
$+\lambda\Big[N_{\frac{\lambda}{2}}\Big(\sum_{s\geq 5}(H_{k}^{s}\chi_{2^{k-1}\leq|.|\leq 2^{k}})*B^{k-s}:k\in{\mathbb{Z}}\Big)(x)\Big]^{\frac{1}{2}}$
@@ 2.2.2. Estimate for $L_{1}$:
$\left|\left\{V_{2}\left(\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}^{d}}H_{j}^{s}*B^{j-s,m}_{u}(x):k\in{\mathbb{Z}}\right)>{c_{1}3^{-d}2^{-\delta_{1}s}\alpha}\right\}\right|,$
for a fixed $u\in\left\{0,\frac{1}{3},\frac{2}{3}\right\}^{d}$. The advantage of the above expression is that $\operatorname{supp}(H_{j}^{s}*B^{j-s}_{u})\subset\widetilde{K}_{j,m}$, where $\widetilde{K}_{j,m}\in\mathcal{D}_{u}$ and any such two $\widetilde{K}_{j,m}$ and $\widetilde{K}_{j^{\prime},m^{\prime}}$ are either disjoint or one is contained in another. We will now discard the points where there is a large interaction on the scales $j$. This leaves us with points having a controlled overlap in the scales with respect to the parameter $s$. To do so, we initially set
$\mathcal{C}_{0}^{u}$ $=\{(j,m):\exists Q\in\mathcal{Q},|Q|=2^{(j-s)d},\;Q\subseteq K_{j,m}\;\text{and}\;\widetilde{K}_{j,m}\in\mathcal{D}_{u}\},\;\text{and}$
$\mathcal{X}^{u}$ $=\bigcup\limits_{(j,m)\in\mathcal{C}^{u}_{0}}\widetilde{K}_{j,m}.$
@@
We note that for any $(j,m)\in\mathcal{C}^{u}_{0}$, there exists atleast one cube $Q_{j.m}\in\mathcal{Q}$ with $|Q_{j,m}|=2^{(j-s)d}$ and $Q_{j,m}\subset K_{j,m}$. Thus, by Chebyshev’s inequality, we have
@@
$\sum_{j\geq k}\sum_{m\in{\mathbb{Z}}}H_{j}^{s}*B^{j-s,m}_{u}(x)=\sum_{(j,m)\in\mathcal{C}^{u},\;j\geq k}H_{j}^{s}*B^{j-s,m}_{u}(x).$
The advantage of working with the collection $\mathcal{C}^{u}$ is that for any fixed $(j,m)\in\mathcal{C}^{u}$, there are atmost $2^{s(d+\epsilon_{0})}$ indices $(j^{\prime},m^{\prime})\in\mathcal{C}^{u}$ such that $\widetilde{K}_{j,m}\subset\widetilde{K}_{j^{\prime},m^{\prime}}$. Based on this observation, we subdivide the collection $\mathcal{C}^{u}$ as follows. For any $(j,m)\in\mathcal{C}^{u}$, we say $(j,m)\in\mathcal{C}^{u}_{1}$ if there is no other index $(j^{\prime},m^{\prime})\in\mathcal{C}^{u}$ such that $\widetilde{K}_{j^{\prime},m^{\prime}}\subseteq\widetilde{K}_{j,m}$. Inductively, we define $\mathcal{C}^{u}_{l}$ to be the collection of all indices $(j,m)\in\mathcal{C}^{u}\setminus\cup_{i=1}^{l-1}\mathcal{C}^{u}_{i}$ such that there does not exist any index $(j,m)\in\mathcal{C}^{u}\setminus\cup_{i=1}^{l-1}\mathcal{C}^{u}_{i}$ such that $\widetilde{K}_{j^{\prime},m^{\prime}}\subset\widetilde{K}_{j,m}$. By the observation before, the process terminates at the step $l=2^{s(d+\epsilon_{0})}$ and we have $\mathcal{C}^{u}=\bigcup\limits_{l=1}^{2^{s(d+\epsilon_{0})}}\mathcal{C}^{u}_{l}$. Thus, we have that
@@
$\lesssim\frac{2^{2\delta_{1}s}}{\alpha^{2}}\left\|V_{2}\left(\sum_{(j,m)\in\mathcal{C}^{u}}H_{j}^{s}*B^{j-s,m}_{u}:k\in{\mathbb{Z}}\right)\right\|_{2}^{2}$
@@ A.1. Estimates of Seeger:
$\int|(A+B)(x)|^{2}dx$
$=$ $\left(\int_{|B(\cdot)|\leq 2^{\rho s}\alpha}+\int_{|A(\cdot)|>2^{\rho s-1}\alpha}+\int_{|(A+B)(\cdot)|>2^{\rho s-1}\alpha}\right)|(A+B)(x)|^{2}dx$
$=:$ $I+II+III.$
@@
$\leq$ $\frac{1}{|R|}\int_{R}\left|\sum_{2^{j}\geq\ell(R)}\nu_{j}*|B^{j-s}|(x)-\frac{1}{|R|}\int_{R}\sum_{2^{j}\geq\ell(R)}\nu_{j}*|B^{j-s}|(y)dy\right|dx$
$+\frac{1}{|R|}\int_{R}\left|\sum_{2^{j}<\ell(R)}\nu_{j}*|B^{j-s}|(x)\right|dx$
@@ Acknowledgement
Ankit Bhojak would like to thank Michael Lacey and Parasar Mohanty for various helpful conversations. Ankit Bhojak is supported by the Science and Engineering Research Board, Department of Science and Technology, Govt. of India, under the scheme National Post-Doctoral Fellowship, file no. PDF/2023/000708.